# Globalization of perturbative Chern-Simons theory on the moduli space of flat connections in the BV formalism

### v1

@@ 1.2.2. Main result 1
    Under 1-extended smoothness assumption (Definition 5.7), [^13] The assumption is that, for smooth irreducible flat connections $A_{0}$, not only the $L_{\infty}$ algebra on $H_{A_{0}}^{\bullet}$ induced from $\Omega^{\bullet}(M,\mathfrak{g})$ vanishes, but also $L_{\infty}$ automorphisms of $H_{A_{0}}^{\bullet}$ induced by variations of the homotopy transfer data (variations of gauge-fixing) vanish. See Definition 5.7 and Remark 5.8. one can recover $Z$ from $Z^{\mathrm{glob}}$ modulo a $\Delta_{\mathsf{a}}$-exact term, as
@@
The generator $R_{1}$ in (18) is given explicitly as a sum over graphs with one marked edge or leaf, cf. Proposition 4.11. Generators $R(A_{0},\alpha;\mathsf{a})$ in (a) and $R_{\mathrm{glob-pert}}$ in (c) are also given explicitly by (195), (213).
@@ 1.5. Extension of the partition function to a nonhomogeneous form in $A,A^{\prime},g$ – main result 4
Here $F$ is the curvature of $\nabla^{\mathbb{H}}$. The expression in brackets acting on $\check{Z}$ above is a flat superconnection concentrated in de Rham degrees 0,1,2 along $\underline{\mathcal{U}}$. By abuse of terminology, we call it the Gauss-Manin superconnection.
@@ 4.5. Metric dependence of the desynchronized partition function
(158) $\widehat{Z}_{A,A^{\prime};\delta g}(\mathsf{a})=e^{\frac{i}{\hbar}S_{CS}(A_{0})}e^{\frac{\pi i}{4}\psi_{A}}\tau_{A}^{1/2}e^{\frac{i}{\hbar}\,\frac{1}{2}\langle\mathsf{a},\widehat{\Theta}(\mathsf{a})\rangle}\;\exp\sum_{\Gamma}\frac{(-i\hbar)^{l(\Gamma)-1}}{|\mathrm{Aut}(\Gamma)|}\widehat{\Phi}_{\Gamma,A,A^{\prime};\delta g}(\mathsf{a})\\ \in\Omega^{0,0,\bullet}(\mathcal{U}\times\mathrm{Met},\mathrm{Dens}^{\frac{1}{2},\mathrm{formal}}(H_{A}))$
@@ 4.6.1. Connection $\nabla^{\mathrm{Hodge}}$
Furthermore, we will denote by $\nabla^{\mathcal{D}}$ the connection induced by $\nabla^{\mathbb{H}}$ on the bundle $\mathcal{D}$ of formal half-densities on $H_{A}$ over $\underline{\mathcal{U}}$.
@@ 4.6.2. Extended partition function
To elucidate the relationship between $\check{Z}$ and $Z$, one can express (171) in terms of the regular Feynman rules, where an edge is assigned $K$ and a leaf is assigned $i(\mathsf{a})$, by adding extra vertices carrying the form degree along $\underline{\mathcal{U}}$: bivalent black and white vertices and grey univalent vertices. White vertices are assigned $H_{\delta A^{\prime}}$, black vertices are assigned $H_{\delta g}$, grey vertices are assigned $\delta A$. In addition, edges decorated by more than two bivalent vertices vanish automatically. See Figure 9. When sandwiched between $K$ and $K$ (or $K$ and $i$) only a single term in $H_{\delta A^{\prime}}$ and $H_{\delta g}$ survives. Therefore, one can evaluate black vertices with $\lambda_{\delta g}$ and white vertices to $dG\mathrm{ad}^{*}_{\delta A^{\prime}}$. The latter effectively acts by applying $(d^{*})^{-1}$” on one of the internal edges incident to the vertex. See for instance the diagram in Figure 10.
@@ 4.6.4. Partial extensions of $Z$ along $A$, $A^{\prime}$ and $g$ – comparison with previous results
or, equivalently,
(187) $(\nabla^{\mathbb{H},A^{\prime}}-\langle[\delta A],\frac{\partial}{\partial\mathsf{a}}\rangle+\langle p\,\mathrm{ad}_{K\delta A}i(\mathsf{a}),\frac{\partial}{\partial\mathsf{a}}\rangle-i\hbar\Delta_{\mathsf{a}})\widetilde{Z}_{\delta A}=0$
@@ 5.1.3. Naive global partition function
Let $\underline{\varphi}\colon U\subset T\mathcal{M}^{\prime}\to\mathcal{M}^{\prime}$ be the sum-over-trees exponential map (69) induced by the SDR $(i_{A},p_{A},K_{A})$. Fix AA and small $\alpha\in H^{1}_{A}$, and let $B=d_{\alpha}\underline{\varphi}_{A}(\alpha)\colon H^{1}_{A}\rightarrow H^{1}_{\underline{\varphi}_{A}(\alpha)}$. Then, for $\mathsf{a}\in H^{1}_{A}$ small, we have
@@
By taking the derivative of (194) in $\alpha$ at $\alpha=0$ one obtains the following.
@@
The partition function $\underline{Z}$ is horizontal w.r.t. Grothendieck connection modulo a BV-exact term:
(197) $\nabla^{G}\underline{Z}=i\hbar\Delta_{\mathsf{a}}\left(r_{A,A;\delta A}(\mathsf{a})Z_{A,A}(\mathsf{a})\right),$
with $r$ as in Proposition 4.11.
5.3. Extended smoothness assumption
Consider the tree-level expression appearing in the extended partition function (138)
(198) $\Xi=\frac{1}{2}\langle\mathsf{a},\widehat{\Theta}(\mathsf{a})\rangle+\sum_{T}\frac{1}{|\mathrm{Aut}(T)|}\widehat{\Phi}_{T,A,A^{\prime}}(\mathsf{a})\in\Omega^{0,\bullet}(\mathcal{U},\mathrm{Sym}(H_{A}[1])^{*})$
with $T$ running over binary rooted trees up to isomorphism; the notations are as in (138).
Definition 5.7.
We will say that $(A,A^{\prime})\in\mathcal{U}$ satisfies the 1-extended smoothness assumption if the $(0,1)$-form component of $\Xi$ vanishes, i.e., if the sum of trees with one edge marked by $\Lambda$ or one leaf marked by $\mathbb{I}i$ or the root marked by $p\mathbb{P}$ (see (125)) is zero. Note that the $(0,0)$-form component of $\Xi$ vanishes by the usual smoothness assumption on $A$. Furthermore, we will say that $(A,A^{\prime})$ satisfies the fully extended smoothness assumption if $\Xi=0$.
1-extended smoothness assumption will be useful for our applications (reduction of the formal exponential map and of the Grothendieck connection to the moduli space), while the fully extended one is given as a natural refinement.
Remark 5.8.
Interpretation of 1-extended smoothness assumption: a variation of $A^{\prime}$ induces a variation of the Hodge SDR data $(i,p,K)$, which in turn induces an $L_{\infty}$ automorphism of the $L_{\infty}$ algebra structure on $H_{A}^{\bullet}$ (with vanishing operations, as per the usual smoothness assumption). The 1-extended smoothness assumption asks that this $L_{\infty}$ automorphism is trivial as well.
Remark 5.9.
Fully extended smoothness assumption can be interpreted using the construction of Appendix C as follows: we are considering the dg Lie algebra $\Omega^{\bullet}(M,\mathfrak{g})\otimes\Omega^{\bullet}(\mathbb{GF}),d_{A}+\delta_{A^{\prime}},[,]$, with $\mathbb{GF}$ as in (286), and we are asking that the homotopy transfer to $H^{\bullet}_{A}\otimes\Omega^{\bullet}(\mathbb{GF})$ yields an $L_{\infty}$ algebra structure with $l_{1}=\delta_{A^{\prime}}$ and all other operations vanishing.
Proposition 5.10.
Consider a path $(A,A^{\prime}_{t})$ in $\mathcal{U}$. Then, under 1-extended smoothness assumption, the sum-over-trees map $\varphi_{A,A^{\prime}_{t}}$ satisfies
(199) $\frac{d}{dt}\varphi_{A,A^{\prime}_{t}}(\mathsf{a})=d_{\varphi_{A,A^{\prime}_{t}}(\mathsf{a})}\gamma$
for sufficiently small $\mathsf{a}$. Here $\gamma$ is given by the sum over trees with $K$ on the root and either one edge (or the root) marked by $\Lambda_{\dot{A}^{\prime}_{t}}$ or one leaf marked by $\mathbb{I}_{\dot{A}^{\prime}_{t}}i(\mathsf{a})$.
In particular, as $t$ changes, $\varphi_{A,A^{\prime}_{t}}(\mathsf{a})$ changes by a gauge transformation.
Sketch of proof.
We have $\varphi_{A,A^{\prime}_{t}}(\mathsf{a})=A+\delta_{A,A^{\prime}_{t}}(\mathsf{a})$, with $\delta$ being the sum over trees with $K$ on the root. Therefore, the l.h.s. of (199) is the sum over trees with one edge (or the root) marked by $d,\Lambda]$ or one leaf marked by $d\mathbb{I}i$ (we suppress the subscripts, in particular, $d=d_{A}$). As in the proof of Proposition 4.11, using the Stokes’ theorem on the configuration space to move $d$ from the marked edge or leaf to other edges, we obtain a sum of:
  1. (1)
  2. (2)
  3. (3)
    Trees with the root marked by $d\Lambda$.
Contributions of 1 on the edge from (1) cancel out in the sum over graphs by the classical master equation (IHX relation). Contributions of $P$ on the edge from (1) and (2) cancel out: (a) if the subtree between $P$ and the leaves does not contain $\mathbb{I},\Lambda$, such subtrees add up to zero by smoothness; (b) if the subtree does contain $\mathbb{I}$ or $\Lambda$, then such subtrees add up to zero by 1-extended smoothness. The remaining contributions are: $dK$ from (2) and $d\Lambda$ from (3) on the root add up to $d\gamma$; $1$ from (2) yields $\delta_{A,A^{\prime}_{t}}(\mathsf{a}),\gamma]$ (from pairs of trees joined at the root, one tree containing $\Lambda$ or $\mathbb{I}$ and one not). Thus, we obtain
(200) $\frac{d}{dt}\varphi_{A,A^{\prime}_{t}}(\mathsf{a})=d\gamma+[\delta_{A,A^{\prime}_{t}}(\mathsf{a}),\gamma]=d_{\varphi_{A,A^{\prime}_{t}}(\mathsf{a})}\gamma.$
In the remainder of the current subsection we will always be assuming 1-extended smoothness.
Let
(201) $\pi\colon\mathrm{FC}\rightarrow\mathcal{M}$
be the quotient map sending a flat connection to its gauge equivalence class.
Corollary 5.11.
We have that
(202) $\pi\varphi_{A,A^{\prime}}(\mathsf{a})=\colon\underline{\varphi}_{A}(\mathsf{a})$
is independent of $A^{\prime}$ close to $A$ and agrees with the formal exponential map (69) on the moduli space $\mathcal{M}^{\prime}$ induced from synchronized Hodge gauge-fixing $(i_{A,A},p_{A,A},K_{A,A})$.
@@ 5.3. Extended smoothness assumption
Obvious from Proposition 5.10. For the second part, set $A^{\prime}=A$. ∎
Corollary 5.12.
The cohomology comparison map
(203) $\mathfrak{B}_{\varphi_{A,A^{\prime}}(\mathsf{a})\leftarrow A,A^{\prime}}\colon H_{A}\rightarrow H_{\varphi_{A,A^{\prime}}(\mathsf{a})}$
is independent of $A^{\prime}$ and coincides with $d_{\mathsf{a}}\underline{\varphi}_{A}$.
Proof.
Follows from Corollary 5.11 by differentiating (202). ∎
Next, consider the setting of Remark 4.8 and set $A^{\prime}=A$. We have
[…]
with $\widetilde{A}=\varphi_{A,A}(\alpha)$ and $\widetilde{\mathsf{a}}=\mathfrak{B}_{\widetilde{A}\leftarrow A,A}(\mathsf{a}-\alpha)$. Comparing with (71), we have the following.
Corollary 5.13.
The restriction of the partial Grothendieck connection $\widetilde{\nabla}^{G}$ (120) to the diagonal $A^{\prime}=A$ upon reduction to the moduli space $\mathcal{M}^{\prime}$ agrees with the Grothendieck connection $\nabla^{G}$ (71) associated with the synchronized Hodge gauge-fixing. More precisely: the parallel transport of $\widetilde{\nabla}^{G}$ from $(A,A)$ to $(\widetilde{A},A)$ coincides with the parallel transport of $\nabla^{G}$ from $[A]$ to $\widetilde{A}$], with $\widetilde{A}=\varphi_{A,A}(\alpha)$ and $\alpha\in H^{1}_{A}$ small.
Corollary 5.14.
The curvature of the connections $\nabla^{\mathrm{Harm}}$ and $\nabla^{\mathbb{H}}$ in the bundle of harmonic forms and the cohomology bundle over $\mathcal{U}\subset\mathrm{FC}^{\prime}\times\mathrm{FC}^{\prime}$ vanishes when restricted to $\delta A$ harmonic (with $\delta A^{\prime}$ arbitrary).
Proof.
The curvature (85) with $\delta A$ harmonic coincides with $O(\delta A\delta A^{\prime})$ term in the contribution of the cubic corolla graph in $\Xi$ (198), evaluated on $\mathsf{a}+[\delta A]$, which vanishes by 1-extended smoothness assumption. ∎
Definition 5.15.
We will say that $(A,A^{\prime},g)\in\underline{\mathcal{U}}$ satisfies metric-extended smoothness if it satisfies the following two properties:
  1. (a)
    1-extended smoothness in the sense of Definition 5.7;
  2. (b)
    the contribution of tree graphs in $r_{\delta g}(\mathsf{a})$ in Theorem 4.17 vanishes.
Metric-extended smoothness implies the obvious generalization of the statements in this section to variations of metric, in particular:
  • The gauge class of $\varphi_{A,A^{\prime}}(\mathsf{a})$ is independent of $g$ and yields $\underline{\varphi}_{A}(\mathsf{a})$. In particular, the latter exponential map on $\mathcal{M}^{\prime}$ does not depend on $g$. Hence, also the Grothendieck connection $\nabla^{G}$ on $\mathcal{M}^{\prime}$, associated with $\underline{\varphi}$, does not depend on $g$.
@@
Throughout this subsection we are assuming 1-extended smoothness.
@@
Step 1. We extend $\underline{Z}$ to a nonhomogeneous form on the moduli space,
(207) $\widetilde{Z}=\underline{Z}+\sum_{p\geq 1}R^{(p)}\in\Omega^{\bullet}(\mathcal{M}^{\prime},\mathrm{Dens}^{\frac{1}{2},\mathrm{formal}}(H_{A}^{\bullet}[1])),$
with $R^{(p)}$ a $p$-form on $\mathcal{M}^{\prime}$ of ghost degree p-p and $R^{(1)}$ being the generator in the r.h.s. of (197), with the extension satisfying
(208) $(\nabla^{G}-i\hbar\Delta_{\mathsf{a}})\widetilde{Z}=0.$
Explicitly, we construct $\widetilde{Z}$ as

### v2

@@ Abstract.
K. W. was supported by the ERC-SyG project, Recursive and Exact New Quantum Theory (ReNewQuantum) with funding from the European Research Council under the European Union‘s Horizon 2020 research and innovation programme, grant agreement no. 810573, and by a Sapere Aude Grant from the Independent Research Fund Denmark, grant ID 10.46540/4251-00015B
@@ 1.2.2. Main result 1
    Under one can recover $Z$ from $Z^{\mathrm{glob}}$ modulo a $\Delta_{\mathsf{a}}$-exact term, as
@@
The generator $R_{1}$ in (18) is given explicitly as a sum over graphs with one marked edge or leaf, cf. Corollary 5.9. Generators $R(A_{0},\alpha;\mathsf{a})$ in (a) and $R_{\mathrm{glob-pert}}$ in (c) are also given explicitly by (211), (231).
@@ 1.5. Extension of the partition function to a nonhomogeneous form in $A,A^{\prime},g$ – main result 4
Here $F$ is the curvature of $\nabla^{\mathbb{H}}$. The expression in brackets acting on $\check{Z}$ above is a flat superconnection concentrated in de Rham degrees 0,1,2 along $\underline{\mathcal{U}}$. By abuse of terminology, we call it the Gauss-Manin superconnection. [^20] In [CMW19], [CMW20] a similar operator was called “Grothendieck operator”. Here, we use the term Gauss-Manin connection since it is a connection on the (half-density bundle of the) cohomology bundle $\mathbb{H}$.
@@ 4.5. Metric dependence of the desynchronized partition function
(158) $\widehat{Z}_{A,A^{\prime};\delta g}(\mathsf{a})=e^{\frac{i}{\hbar}S_{CS}(A)}e^{\frac{\pi i}{4}\psi_{A}}\tau_{A}^{1/2}e^{\frac{i}{\hbar}\,\frac{1}{2}\langle\mathsf{a},\widehat{\Theta}(\mathsf{a})\rangle}\;\exp\sum_{\Gamma}\frac{(-i\hbar)^{l(\Gamma)-1}}{|\mathrm{Aut}(\Gamma)|}\widehat{\Phi}_{\Gamma,A,A^{\prime};\delta g}(\mathsf{a})\\ \in\Omega^{0,0,\bullet}(\mathcal{U}\times\mathrm{Met},\mathrm{Dens}^{\frac{1}{2},\mathrm{formal}}(H_{A}[1]))$
@@ 4.6.1. Connection $\nabla^{\mathrm{Hodge}}$
Furthermore, we will denote by $\nabla^{\mathcal{D}}$ the connection induced by $\nabla^{\mathbb{H}}$ on the bundle $\mathcal{D}$ of formal half-densities on $H_{A}{[1]}$ over $\underline{\mathcal{U}}$.
@@ 4.6.2. Extended partition function
To elucidate the relationship between $\check{Z}$ and $Z$, one can express (171) in terms of the regular Feynman rules, where an edge is assigned $K$ and a leaf is assigned $i(\mathsf{a})$, by adding extra vertices carrying the form degree along $\underline{\mathcal{U}}$: bivalent black and white vertices and gray univalent vertices. White vertices are assigned $H_{\delta A^{\prime}}$, black vertices are assigned $H_{\delta g}$, gray vertices are assigned $\delta A$. In addition, edges decorated by more than two bivalent vertices vanish automatically. See Figure 9. When sandwiched between $K$ and $K$ (or $K$ and $i$) only a single term in $H_{\delta A^{\prime}}$ and $H_{\delta g}$ survives. Therefore, one can evaluate black vertices with $\lambda_{\delta g}$ and white vertices to $dG\mathrm{ad}^{*}_{\delta A^{\prime}}$. The latter effectively acts by applying $(d^{*})^{-1}$” on one of the internal edges incident to the vertex. See for instance the diagram in Figure 10.
@@ 4.6.4. Partial extensions of $Z$ along $A$, $A^{\prime}$ and $g$ – comparison with previous results
which implies
(187) $(\nabla^{\mathbb{H},A^{\prime}}-\langle[\delta A],\frac{\partial}{\partial\mathsf{a}}\rangle+\langle p\,\mathrm{ad}_{K\delta A}i(\mathsf{a}),\frac{\partial}{\partial\mathsf{a}}\rangle+O((\delta A)^{2})-i\hbar\Delta_{\mathsf{a}})\widetilde{Z}_{\delta A}=0$
@@ 5.1.3. Naive global partition function
5.2. A formula for the Grothendieck connection
First we introduce some notations that we will be using in the remainder of this section.
Let $\Xi$ be the one-form component (as a form on $\underline{\mathcal{U}}$) of the tree part of the extended partition function $\check{Z}$ (171):
(194) $\Xi=-\langle[\delta A],\mathsf{a}\rangle+\frac{1}{2}\langle\mathsf{a},\check{\Theta}(\mathsf{a})\rangle+\sum_{T}\frac{1}{|\mathrm{Aut}(T)|}\check{\Phi}_{T}(\mathsf{a})\Big|_{\Omega^{1}(\underline{\mathcal{U}})}\quad\in\Omega^{1}(\underline{\mathcal{U}},\widehat{\mathrm{Sym}}(H_{A}[1])^{*})$
with $T$ running over trivalent trees, with notations as in (171). $\Xi$ splits as a 1-from along $A$ plus a 1-from along $A^{\prime}$ plus a 1-form along $g$:
(195) $\Xi=\Xi_{\delta A}+\Xi_{\delta A^{\prime}}+\Xi_{\delta g}.$
Furthermore, let
(196) $I^{*}\colon\Omega^{\bullet}(\mathcal{U})\rightarrow\Gamma(\mathrm{FC}^{\prime},\wedge^{\bullet}(H^{1}_{A})^{*})$
be the map which restricts a form on $\mathcal{U}\subset\mathrm{FC}^{\prime}\times\mathrm{FC}^{\prime}$ to the diagonal $A^{\prime}=A$ and evaluates it on $(A,A)$-harmonic tangent vectors $\delta A^{\prime}=\delta A$.
Then, given a form $\omega$ on $\mathcal{U}$ valued in the bundle $\mathcal{D}=\mathrm{Dens}^{\frac{1}{2},\mathrm{formal}}(H_{A}[1])$ and invariant under gauge transformations, we can construct a $\mathcal{D}$-valued form on $\mathcal{M}^{\prime}$ by evaluating $\omega$ on the diagonal $A^{\prime}=A$ and harmonic tangent vectors $\delta A^{\prime}=\delta A$, and passing to gauge-equivalence classes. I.e., we have a map
(197) $\pi_{*}I^{*}\colon\Omega^{\bullet}(\mathcal{U},\mathcal{D})^{\mathrm{Gauge}}\rightarrow\Omega^{\bullet}(\mathcal{M}^{\prime},\mathcal{D}).$
Here $\pi\colon\mathrm{FC}\rightarrow\mathcal{M}$ is the projection to gauge-equivalence classes. The superscipt “Gauge” means “invariants under the action of the gauge group.”
Denote
[…]
– it is the term $\langle[\delta A],\mathsf{a}\rangle$ plus the sum of trees with one white binary vertex $H_{\delta A^{\prime}=i[\delta A]}$, in terms of Remark 4.22.
Proposition 5.5.
Grothendieck connection associated to the formal exponential map $\underline{\varphi}$ on $\mathcal{M}^{\prime}$, seen as an operator on $\Omega^{\bullet}(\mathcal{M}^{\prime},\mathcal{D})$, can be written as
(199) $\nabla^{G}=\nabla^{\mathbb{H}}_{\mathcal{M}}+\{\xi,-\}+\Delta_{\mathsf{a}}\xi.$
Here $\nabla^{\mathbb{H}}_{\mathcal{M}}=\pi_{*}I^{*}\nabla^{\mathbb{H}}$ is the connection on the bundle $\mathcal{D}\rightarrow\mathcal{M}^{\prime}$ obtained by restricting the connection in the cohomology bundle $\nabla^{\mathbb{H}}$ over $\mathcal{U}$ (induced by $\nabla^{\mathrm{Harm}}$ of Section 2.5.3) to the diagonal $A^{\prime}=A$ and harmonic tangent vectors, and passing to gauge-equivalence classes.
For the proof of Proposition 5.5 we will need the following lemma.
Lemma 5.6.
Consider a path $(A,A^{\prime}_{t})$ in $\mathcal{U}$ passing through $(A,A^{\prime})$ at $t=0$. Then the sum-over-trees map $\varphi_{A,A^{\prime}_{t}}\colon H^{1}_{A}\supset U\rightarrow\mathrm{FC}^{\prime}$ satisfies
(200) $\frac{d}{dt}\Big|_{t=0}\varphi_{A,A^{\prime}_{t}}(\mathsf{a})=d_{\varphi_{A,A^{\prime}}(\mathsf{a})}\gamma+(d_{\mathsf{a}}\varphi_{A,A^{\prime}})(\beta)$
for sufficiently small $\mathsf{a}$. Here:
  • $\gamma$ is given by the sum over trees with $K$ on the root and either one edge (or the root) marked by $\Lambda_{\dot{A}^{\prime}}$ or one leaf marked by $\mathbb{I}_{\dot{A}^{\prime}}i(\mathsf{a})$. One also allows for $\gamma$ a special graph consisting of a single marked leaf with value $\mathbb{I}_{\dot{A}^{\prime}}i(\mathsf{a})$.
  • $\beta=\{\Xi_{\delta A^{\prime}=\dot{A}^{\prime}},-\}$ is given by the sum over trees with $p$ on the root, with one edge marked by $\Lambda_{\dot{A}^{\prime}}$ or one leaf marked by $\mathbb{I}_{\dot{A}^{\prime}}i(\mathsf{a})$ or the root marked by $p\mathbb{P}_{\dot{A}^{\prime}}$.
Note that the first term in the r.h.s. of (200) is an infinitesimal gauge transformation.
Sketch of proof.
We have $\varphi_{A,A^{\prime}_{t}}(\mathsf{a})=A+\delta_{A,A^{\prime}_{t}}(\mathsf{a})$, with $\delta$ being the sum over trees with $K$ on the root. Therefore, the l.h.s. of (200) is the sum over trees with one edge (or the root) marked by $d,\Lambda]+\mathbb{I}P+P\mathbb{P}$ or one leaf marked by $d\mathbb{I}i$ (we suppress the subscripts, in particular, $d=d_{A}$). As in the proof of Proposition 4.11, using the Stokes’ theorem on the configuration space to move $d$ from the marked edge or leaf to other edges, we obtain a sum of:
  1. (1)
  2. (2)
  3. (3)
    Trees with the root marked by $d\Lambda$.
  4. (4)
    Trees with one edge (or root) decorated by $\mathbb{I}P$.
  5. (5)
    Trees with one edge (or root) decorated by $P\mathbb{P}$.
Contributions of 1 on the edge from (1) cancel out in the sum over graphs by the classical master equation (IHX relation). Contributions of $P$ on the edge from (1) and (2), together with (4) and (5), add up to the second term in the r.h.s. of (200): (a) if the subtree between $P$ and the leaves does not contain $\mathbb{I},\Lambda$, such subtrees add up to zero by smoothness, and the same applies to (4); (b) if the subtree does contain $\mathbb{I}$ or $\Lambda$, then such subtrees, as well as the ones coming from with (5), add up to $\beta$. The remaining contributions are: $dK$ from (2) and $d\Lambda$ from (3) on the root add up to $d\gamma$; $1$ from (2) yields $\delta_{A,A^{\prime}}(\mathsf{a}),\gamma]$ (from pairs of trees joined at the root, one tree containing $\Lambda$ or $\mathbb{I}$ and one not). Thus, we obtain
(201) $\frac{d}{dt}\Big|_{t=0}\varphi_{A,A^{\prime}_{t}}(\mathsf{a})=d\gamma+[\delta_{A,A^{\prime}}(\mathsf{a}),\gamma]+(d_{\mathsf{a}}\varphi_{A,A^{\prime}})(\beta)=d_{\varphi_{A,A^{\prime}}(\mathsf{a})}\gamma+(d_{\mathsf{a}}\varphi_{A,A^{\prime}})(\beta).$
Proof of Proposition 5.5.
Fix $A\in\mathrm{FC}^{\prime}$ and $\mathsf{a}\in H^{1}_{A}$ sufficiently small. Consider the first-order deformation $A\rightarrow\widetilde{A}=A+\epsilon i(\alpha)$ with $\alpha\in H^{1}_{A}$. Then the infinitesimal parallel transport of $\nabla^{G}$, seen as a fiber bundle connection on a neighborhood of the zero-section in $T\mathcal{M}^{\prime}$ (Definition 2.21), maps $\mathsf{a}\rightarrow\widetilde{\mathsf{a}}$ where $\widetilde{\mathsf{a}}$ is determined by
[…]
We will write $\widetilde{\mathsf{a}}=B(\mathsf{a}+\epsilon\mathsf{b})$ (all computations are $\bmod\,\epsilon^{2}$), with $B=\mathfrak{B}^{1}_{\widetilde{A}\leftarrow A;\widetilde{A}}\colon H^{1}_{A}\rightarrow H^{1}_{\widetilde{A}}$ the cohomology comparison map. [^52] Note that in the first order in $\epsilon$ it does not matter whether we choose $A^{\prime}=A$ or $A^{\prime}=\widetilde{A}$ in the cohomology comparison map $\mathfrak{B}_{\widetilde{A}\leftarrow A;A^{\prime}}$. Equation (202) reads
[…]
(205) $=[\varphi_{A,\widetilde{A}}(\mathsf{a})+\epsilon(d_{\mathsf{a}}\varphi_{A,A})(\alpha+\mathsf{b})]$
[…]
(207) $=[\varphi_{A,A}(\mathsf{a})+\epsilon(d_{\mathsf{a}}\varphi_{A,A})(\alpha+\mathsf{b}+\beta)]$
Here $\beta,\gamma$ are as in Lemma 5.6 with $A^{\prime}=A$ and $\dot{A}^{\prime}=i(\alpha)$. Thus, (202) holds if the $O(\epsilon)$ term in the last line vanishes, i.e., if $\mathsf{b}=-\alpha-\beta$ and thus
(208) $\widetilde{\mathsf{a}}=B(\mathsf{a}-\epsilon(\alpha+\beta)).$
Therefore, as a fiber bundle connection on (a neighborhood of the zero-section in) $T\mathcal{M}^{\prime}$,
(209) $\nabla^{G}=\nabla^{\mathbb{H}}_{\mathcal{M}}+\langle[\delta A],\frac{\partial}{\partial\mathsf{a}}\rangle+\{\pi_{*}I^{*}\Xi_{\delta A^{\prime}=i[\delta A]},-\}.$
Its action on half-densities is given by (199). ∎
Remark 5.7.
When we consider $\nabla^{G}$ as a fiber bundle connection on $T\mathcal{M}^{\prime}$, $\mathsf{a}=\mathsf{a}^{1}$ is always an element of $H^{1}_{A}$; second and third terms in (209) give a 1-form on $\mathcal{M}^{\prime}$ valued in formal vector fields on $H^{1}_{A}$. On the other hand, when we allow $\nabla^{G}$ to act on half-densities on $H^{\bullet}_{A}[1]$, we have the cotangent lift of that vector field, $\{\xi,-\}$ (which involves $\mathsf{a}^{2}$ – the component $H^{2}_{A}$ of $\mathsf{a}$) appearing in (199). The term $\Delta_{\mathsf{a}}\xi$ arises from the natural action of a hamiltonian vector field on $H^{\bullet}_{A}[1]$ on a half-density.
@@ 5.2. A formula for the Grothendieck connection
Let $\underline{\varphi}\colon U\subset T\mathcal{M}^{\prime}\to\mathcal{M}^{\prime}$ be the sum-over-trees exponential map (69) induced by the SDR $(i_{A},p_{A},K_{A})$. Fix $A$ and small $\alpha\in H^{1}_{A}$, and let $B=d_{\alpha}\underline{\varphi}_{A}(\alpha)\colon H^{1}_{A}\rightarrow H^{1}_{\underline{\varphi}_{A}(\alpha)}$. Then, for $\mathsf{a}\in H^{1}_{A}$ small, we have
@@
Partition function $\underline{Z}$ is horizontal w.r.t. Grothendieck connection modulo a BV-exact term:
(213) $\nabla^{G}\underline{Z}=i\hbar\Delta_{\mathsf{a}}(\underline{r}\,\underline{Z}),$
where $\underline{r}$ is $r_{A,A;i[\delta A]}$ of Proposition 4.11 with trees removed. I.e., $\underline{r}$ is the sum of connected Feynman graphs with $\geq 1$ loops, with one edge marked by $\Lambda_{i[\delta A]}$ or one leaf marked by $\mathbb{I}_{i[\delta A]}i(\mathsf{a})$.
@@
Taking the derivative of (210) in $\alpha$ at $\alpha=0$ one obtains the following:
(214) $\nabla^{G,\mathrm{triv}}\underline{Z}=i\hbar\Delta_{\mathsf{a}}(\pi_{*}I^{*}(r)\,\underline{Z}),$
where $r$ is as in Proposition 4.11 and $\nabla^{G,\mathrm{triv}}\colon=\nabla^{\mathbb{H}}_{\mathcal{M}}+\langle[\delta A],\frac{\partial}{\partial\mathsf{a}}\rangle$. Adding $\Delta_{\mathsf{a}}(\pi_{*}I^{*}\Xi_{\delta A^{\prime}}\,\underline{Z})$ to both sides and using Proposition 5.5, we obtain (213). ∎
@@
For the proof (and for the proof of Proposition 5.19 below) we will need the following lemma.
Lemma 5.11.
  1. (a)
    Let
    (217) $\overline{Z}\colon=\pi_{*}I^{*}(e^{-\frac{i}{\hbar}\Xi}\check{Z})\quad\in\Omega^{\bullet}(\mathrm{Met}\times\mathcal{M}^{\prime},\mathcal{D})$
    and
    (218) $\overline{Z}^{\mathrm{ren}}\colon=e^{\frac{i}{\hbar}c(\hbar)\frac{S_{\mathrm{grav}}(g,\phi)}{2\pi}}\overline{Z}.$
    Then $\overline{Z}^{\mathrm{ren}}$ satisfies
    (219) $(\delta_{g}+\nabla^{G}+\{\mu,-\}+\Delta_{\mathsf{a}}\mu-i\hbar\Delta_{\mathsf{a}})\overline{Z}^{\mathrm{ren}}=0.$
    Here $\mu=\pi_{*}I^{*}\Xi_{\delta g}\in\Omega^{1,0}(\mathrm{Met}\times\mathcal{M}^{\prime},\widehat{\mathrm{Sym}}(H_{A}[1])^{*})$.
  2. (b)
    Let
    (220) $\widetilde{Z}\colon=\overline{Z}|_{g\mathrm{\;fixed,\;}\delta g=0}\;\in\Omega^{\bullet}(\mathcal{M}^{\prime},\mathcal{D})$
    – the restriction of $\overline{Z}$ to a $g$-fixed slice of $\mathrm{Met}\times\mathcal{M}^{\prime}$. [^54] Note that $\widetilde{Z}=\underline{Z}+\underline{r}\,\underline{Z}+\cdots$ with $\underline{r}\,\underline{Z}$ the generator in the r.h.s. of (213) and $\cdots$ being a sum of forms of degree $\geq 2$ on $\mathcal{M}^{\prime}$. Then $\widetilde{Z}$ satisfies
    (221) $(\nabla^{G}-i\hbar\Delta_{\mathsf{a}})\widetilde{Z}=0.$
@@
Denote $\xi^{\mathrm{tot}}\colon=\pi_{*}I^{*}\Xi=\xi+\mu$. The dQME (175) implies